# Riemannian Optimization for Distance Geometry: A Study of Convergence, Robustness, and Incoherence

### v1

@@ 7.1 A Riemannian Approach to Matrix Completion
7.2 Euclidean Distance Geometry Algorithms
To solve the EDG problem, various algorithms have been developed. Among them, one prominent family of algorithms is based on semi-definite programming (SDP), which leverages the connection between squared distance matrices and Gram matrices. To provide a concrete example of this approach, we briefly outline the method proposed in [67]. Consider the matrix $\bm{V}\in\mathbb{R}^{n\times(n-1)}$, whose columns form an orthonormal basis for the space $\{\bm{z}\in\mathbb{R}^{n}:\bm{z}^{\top}\bm{1}=0\}$. The operator $\mathcal{K}$ is defined as:
@@ 7.2 Euclidean Distance Geometry Algorithms
We refer the reader to [67] for theoretical and numerical aspects of the above optimization program. Given that standard SDP formulations can be computationally intensive, distributed and divide-and-conquer methods have also been explored. For additional SDP-based formulations of the EDG problem and their applications to molecular conformation and sensor network localization, we refer the reader to [6, 68, 69, 70, 71, 56].
In the context of protein structure determination, various algorithmic approaches to EDG have been developed. One notable example is the EMBED algorithm[72, 73, 74], which comprises three main steps[75]. The first step, known as bound smoothing, involves generating lower and upper bounds for all distances by extrapolating from the available limits of known distances. The second step is the embed step, where distances are sampled from these bounds to form a full distance matrix from which an initial estimate of the protein structure is obtained. The final step involves refining this initial structure by minimizing an energy function using non-convex optimization methods. Another approach to structure prediction is the discretizable molecular distance geometry framework, which can be formulated as a search in a discrete space followed by a Branch-and-Prune method [76, 77].
Another category of approaches to the EDG problem involves initially estimating a smaller portion of the point cloud and then using this initial estimate to incrementally reconstruct the rest of of the structure. These methods are referred to as geometric build-up algorithms[78, 79, 80]. The algorithm proposed in [81] addresses the molecular conformation problem by adopting a divide-and-conquer strategy, where a sequence of smaller optimization problems is solved instead of solving a single global optimization problem.
@@
We note that the above discussion does not comprehensively cover all EDG algorithms, and we refer readers to [88, 20] for a more detailed overview.
7.2.1 Related Geometric Approaches to EDG
The main perspective taken in this paper is in line with low-rank matrix completion approach, albeit not one that employs the trace heuristic seen in [28, 6, 89]. This work is more in line with non-convex approaches based on optimizing over a Riemannian manifold [32, 90], and extends the Riemannian approach of [51] to the EDG basis case.
A recent work in [30] adopts a similar approach to us and considers solving the EDG problem through Riemannian methods as well. In this work, the authors use a Riemannian conjugate method paired with an inexact line search method to minimize the following s-stress objective function:
@@ 7.2.1 Related Geometric Approaches to EDG
where $g$ is the map defined by (2), $\bm{W}$ is a weight matrix to model noisy entries, and $\odot$ is the Hadamard product, and $\mathcal{P}_{\Omega}$ is defined as in (6). The analysis in [30] centers around the minimization of the s-stress function in (28) using a generalization of a Hager-Zhang line search method to a Riemannian quotient manifold. The main result in this work is that there exists an attractive basin for (28) that, with high probability, gives linear convergence to the ground truth provided an initialization in the basin. This result requires a Bernoulli sample complexity $p>C\frac{(\nu r)^{3}\log{n}}{n}$ , where $\nu$ is the incoherence of the ground truth matrix and $r$ is the rank. In contrast, our method also shows linear convergence in a local neighborhood and describes a strong initialization candidate for the noiseless EDG recovery problem with provable high probability guarantees. We also provide robustness analysis for an EDG problem perturbed by noise, and provide provable guarantees as well.
@@ 8 Numerical Results
All of the following experiments were conducted in MATLAB. The code used for the following experiments can be found in the GitHub repository at https://github.com/chandlersmith2/Riemannian_EDG.
@@ 8.1 Synthetic Data Experiment
Sphere (3D, $n=1002$) 3.38e-07 4.61e-07 6.12e-07 1.48e-06 8.40e-03 6.81e-01
Cow (3D, $n=2601$) 4.41e-07 5.24e-07 6.04e-07 9.14e-07 2.47e-04 5.71e-03
Swiss Roll (3D, $n=2048$) 3.85e-07 4.70e-07 5.81e-07 9.47e-07 1.56e-06 6.40e-02
[…]
Figure 4: Reconstruction of the synthetic datasets referenced in Table 2. From left to right, the Bernoulli parameter is 0.030.030.03, 0.020.020.02, and 0.010.010.01.
@@ 8.2 Comparison to existing methods
We provide an additional experiment to compare the efficacy of Algorithm 1 to another provably convergent non-convex EDG algorithm[35]. Let $r\in[2,10]$, and consider $n=100$ points sampled from $\mathrm{Unif}(S^{r-1})$, the uniform distribution on the sphere embedded in $r$ dimensions. As the number of degrees of freedom in a $r$ $n\times n$ matrix is $nr-\frac{r(r-1)}{2}$, define the oversampling ratio $\rho$ as
@@
as $\mathbb{E}[|\Omega|]=pL$ for Bernoulli random sampling with parameter $p$. In Figure 5, we compare the oversampling ratio versus the dimension of the sphere in a transition plot. Black indicates complete failure, classified as a relative Gram matrix error larger than $10^{-3}$, and white indicates success. Each of these squares were run for 100 trials using Algorithm 1 and the algorithm in [35]. Algorithm 1 was initialized with 10 iterations of the algorithm in [28], and [35] was initialized using a least-squares methodology described in that paper.
Algorithm 1
[35]
Figure 5: Oversampling ratio $\rho$ versus dimension $r$ for 100100100 points on the uniform distribution on $S^{r-1}$. Each parameter was tested 100 times.
As these experiments indicate, performance between [35] and Algorithm 1 is comparable, with Algorithm 1 performing slightly better overall, but most noticeably in the higher rank regime.
@@ 8.3 Experiments on Noisy Distance Measurements
Figure 6: Oversampling ratio $\rho$ versus noise level $10^{\gamma}$ for 100100100 points drawn i.i.d. from $\mathrm{Unif}(S^{2})$. Each parameter was tested 500 times.
@@ 9 Conclusion and Future Work
In this work, we proposed a novel Riemannian gradient descent approach for solving the EDG problem using a matrix completion approach on the manifold of $r$ matrices in Algorithm 1. In a local neighborhood, we proved that Algorithm 1 exhibits linear convergence with high probability. To the authors’ knowledge, this is the first work to provide initialization guarantees for a non-convex approach to the EDG problem. The convergence analysis of Algorithm 1 was predicated on a statistical understanding of coupled terms in a random operator, and required novel analysis to the matrix completion literature to our knowledge. For our method, we provided numerical results to underline its efficacy, and Algorithm 1 performs comparably to other state-of-the-art non-convex methods. Additionally, we provided robustness analysis and corresponding convergence guarantees. Finally, we provided a novel interpretation of incoherence in the EDG setting, highlighting potential areas for development of non-uniform sampling methods in this field. This is a primary avenue of future interest, as improving the sample complexity through geometrically-optimal sampling schemes would represent a noteworthy development in the EDG literature.
10 Acknowledgment
Abiy Tasissa and Chandler Smith acknowledge partial support from the National Science Foundation through grant DMS-2208392. HanQin Cai acknowledges partial support from the National Science Foundation through grant DMS-2304489.
@@ Appendix A Properties of the dual bases and Concentration Inequalities
This section of the appendix details technical results about the specific dual bases, $\{\bm{w}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$ and $\{\bm{v}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$. These are needed to prove various technical lemmas throughout the work, but are particularly important in the proof of Theorem 5.3. Additionally, we provide the non-commutative and scalar Bernstein inequalities, as well as the Hanson-Wright inequality and the Davis-Kahan Theorem, all of which are leveraged throughout this work.
@@
$\mathbb{P}\left(\left\|\sum_{i=1}^{m}\bm{X}_{i}\right\|>t\right)\leq 2n\exp\left(-\frac{3t^{2}}{8\sigma^{2}}\right).$ (29)
and if $t>\frac{\sigma^{2}}{c}$
@@
$\delta=\inf\left\{\left|\lambda-\hat{\lambda}\right|:\lambda\in[\lambda_{s},\lambda_{r}],\hat{\lambda}\in\left(-\infty,\hat{\lambda}_{s+1}\right)\cup\left(\hat{\lambda}_{r-1},\infty\right)\right\}$
@@
In particular, for $r$ matrices $\bm{X},\hat{\bm{\bm{X}}}\succeq\bm{0}$ with eigenvectors corresponding to non-zero eigenvalues forming the columns of $\bm{V},\hat{\bm{V}}$, $\delta=\lambda_{r}$ and
@@
One result that will be used throughout this work is a technique for constructing eigenvalue bounds through a vectorization technique. This result is as follows.
@@
Lemma A.6 ($\lambda_{\mathrm{max}}(\tilde{\bm{H}})$ bound).
Let $\tilde{\bm{H}}=[\langle\mathcal{P}_{U}\bm{w}_{\bm{\alpha}},\mathcal{P}_{U}\bm{w}_{\bm{\beta}}\rangle]\in\mathbb{R}^{L\times L}$, where $U$ is the row/column space of the true solution $\bm{X}=\bm{U}\bm{D}\bm{U}^{\top}$, which is $r$, and where $\mathcal{P}_{U}$ is the projection operator onto $U$. It follows that
@@
[…]
@@
$\sum_{\bm{\alpha}\in\mathbb{I}}\langle\bm{X}\mathcal{P}_{U},\mathcal{P}_{U}\bm{w}_{\bm{\alpha}}\rangle^{2}\leq\max_{\|\bm{X}\|_{\mathrm{F}}=1}\sum_{\bm{\alpha}\in\mathbb{I}}\langle\bm{X}\mathcal{P}_{U},\mathcal{P}_{U}\bm{w}_{\bm{\alpha}}\rangle^{2}\leq\lambda_{\mathrm{max}}(\tilde{\bm{H}}).$
@@
The second statement follows from Lemma A.5 and the fact that $\mathcal{P}_{U}$ is an orthogonal projection operator. This concludes the proof. ∎
@@
Let $\bm{X}_{l}=\bm{U}_{l}\bm{D}_{l}\bm{U}_{l}^{\top}$ be a $r$ matrix and $\mathbb{T}_{l}$ be the tangent space of $\mathcal{N}_{r}$ at $\bm{X}_{l}$. Let $\bm{X}=\bm{U}\bm{D}\bm{U}^{\top}$ be another $r$ matrix, and $\mathbb{T}$ be the corresponding tangent space. Then
$\|\bm{U}_{l}\bm{U}_{l}^{\top}-\bm{U}\bm{U}^{\top}\|\leq\frac{\|\bm{X}_{l}-\bm{X}\|_{\mathrm{F}}}{\sigma_{r}(\bm{X})},\qquad\qquad\|\bm{U}_{l}\bm{U}_{l}^{\top}-\bm{U}\bm{U}^{\top}\|_{\mathrm{F}}\leq\frac{\sqrt{2}\|\bm{X}_{l}-\bm{X}\|_{\mathrm{F}}}{\sigma_{r}(\bm{X})}$
@@ Appendix B Restricted Isometry Results
$(\mathcal{S}_{\Omega}(\bm{X}))_{\bm{\alpha}}$ =$=\begin{cases}\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle&\bm{\alpha}\in\Omega\\ 0&\bm{\alpha}\not\in\Omega\end{cases}$ (31)
@@
[…]
@@
$(H^{-1}_{\mathrm{offdiag}})_{\bm{\alpha}\bm{\beta}}=\begin{cases}H^{\alpha\beta}&\bm{\alpha}\neq\bm{\beta}\\ 0&\bm{\alpha}=\bm{\beta}.\end{cases}$ (33)
@@
$\leq\exp\left(\lambda^{2}p\|\bm{A}\|_{\mathrm{F}}^{2}\|\bm{B}\|_{\mathrm{F}}^{2}\right).$
@@
Let $\hat{\mathcal{F}_{\Omega}}$ be either $\mathcal{P}_{\mathbb{T}}\mathcal{F}_{\Omega}\mathcal{P}_{\mathbb{T}}$ or $\mathcal{F}_{\Omega}$, and let ^$\hat{\mathcal{F}}_{\mathbb{I}}$ be $\mathcal{P}_{\mathbb{T}}\mathcal{F}_{\mathbb{I}}\mathcal{P}_{\mathbb{T}}$ or $\mathcal{F}_{\mathbb{I}}$, respectively. Let $\hat{\bm{w}}_{\bm{\alpha}}$ be either $\mathcal{P}_{\mathbb{T}}\bm{w}_{\bm{\alpha}}$ or $\bm{w}_{\bm{\alpha}}$, respectively. Let $\hat{\bm{H}}=[\langle\hat{\bm{w}}_{\bm{\alpha}},\hat{\bm{w}}_{\bm{\beta}}\rangle]\in\mathbb{R}^{L\times L}$ be the corresponding correlation matrix. For a ground truth rank $r$ $\nu$-incoherent matrix $\bm{X}$ with tangent space $\mathbb{T}$ on $\mathcal{N}_{r}$, we have that for any $\beta>1$, and with probability at least $1-2n^{1-\beta}$, and for $p\geq\frac{4\beta\log n}{3n}$, that
^$\|\hat{\mathcal{F}}_{\Omega}-p\hat{\mathcal{F}}_{\mathbb{I}}\|\leq p\sqrt{\frac{8\beta\left(\max_{\bm{\alpha}}\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\right)\lambda_{\max}(\hat{\bm{H}})\log n}{3p}}.$
@@
[…]
@@
where the last inequality follows from Assumption 5.1. Next, we seek to bound the variance term, $\sigma^{2}=\left\|\sum_{\bm{\alpha}\in\mathbb{I}}\mathbb{E}\left[\bm{S}_{\bm{\alpha}}^{2}\right]\right\|$. To see this, first notice that
@@
$\leq p\left(\max_{\bm{\alpha}}\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\right)\lambda_{\max}(\hat{\bm{H}}).$
@@
For a ground truth $r$, $\nu$-incoherent matrix $\bm{X}$ with tangent space $\mathbb{T}$ on $\mathcal{N}_{r}$, we have that for any $\beta>1$ that if $p\geq\frac{4}{3}\frac{\beta\log{n}}{n}$, with probability at least $1-2n^{1-\beta}$
@@
We will prove this result using Theorem A.1. First, notice that
@@
where the first inequality follows from dropping the negative term, and the third inequality follows from Assumption 5.1 and $\|\bm{w}_{\bm{\alpha}}\|_{\mathrm{F}}=2$. Next, we note that
@@
where the first inequality follows from Assumption 5.1, the second inequality comes from Lemma A.5, and the final line comes from Lemma A.8. Next, notice that
@@
where the first inequality comes from $\|\bm{w}_{\bm{\alpha}}\|_{F}^{2}=4$, the second inequality comes from Lemma A.5, and the final line comes from $\lambda_{\max}(\tilde{\bm{H}})\leq\nu r$ in Lemma A.6. Taking $\sigma^{2}=\max\{\sigma_{1}^{2},\sigma_{2}^{2}\}=4p\nu r$, we get that for $t=\sqrt{\frac{32p\nu r\beta\log{n}}{3}}$ and $p\geq\frac{4}{3}\frac{\beta\log{n}}{{n}}$,
@@
$\left|\mathcal{S}_{\Omega}^{\top}(\hat{\bm{Y}})\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\hat{\bm{Y}})-\mathbb{E}\left[\mathcal{S}_{\Omega}^{\top}(\hat{\bm{Y}})\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\hat{\bm{Y}})\right]\right|$ $\leq Cp\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\beta\log{n}$

### v2

@@ 7.1 A Riemannian Approach to Matrix Completion
7.2 Euclidean Distance Matrix Completion Algorithms
To solve the EDMC problem, various algorithms have been developed. Among them, one prominent family of algorithms is based on semi-definite programming (SDP), which leverages the connection between squared distance matrices and Gram matrices. To provide a concrete example of this approach, we briefly outline the method proposed in [70]. Consider the matrix $\bm{V}\in\mathbb{R}^{n\times(n-1)}$, whose columns form an orthonormal basis for the space $\{\bm{z}\in\mathbb{R}^{n}:\bm{z}^{\top}\bm{1}=0\}$. The operator $\mathcal{K}$ is defined as:
@@ 7.2 Euclidean Distance Matrix Completion Algorithms
We refer the reader to [70] for theoretical and numerical aspects of the above optimization program. Given that standard SDP formulations can be computationally intensive, distributed and divide-and-conquer methods have also been explored. For additional SDP-based formulations of the EDMC problem and their applications to molecular conformation and sensor network localization, we refer the reader to [6, 71, 72, 73, 74, 59].
In the context of protein structure determination, various algorithmic approaches to EDMC have been developed. One notable example is the EMBED algorithm[75, 76, 77], which comprises three main steps[78]. The first step, known as bound smoothing, involves generating lower and upper bounds for all distances by extrapolating from the available limits of known distances. The second step is the embed step, where distances are sampled from these bounds to form a full distance matrix from which an initial estimate of the protein structure is obtained. The final step involves refining this initial structure by minimizing an energy function using non-convex optimization methods. Another approach to structure prediction is the discretizable molecular distance geometry framework, which can be formulated as a search in a discrete space followed by a Branch-and-Prune method [79, 80].
Another category of approaches to the EDMC problem involves initially estimating a smaller portion of the point cloud and then using this initial estimate to incrementally reconstruct the rest of of the structure. These methods are referred to as geometric build-up algorithms[81, 82, 83]. The algorithm proposed in [84] addresses the molecular conformation problem by adopting a divide-and-conquer strategy, where a sequence of smaller optimization problems is solved instead of solving a single global optimization problem.
@@
We note that the above discussion does not comprehensively cover all EDMC algorithms, and we refer readers to [91, 20] for a more detailed overview.
7.2.1 Related Geometric Approaches to EDMC
The main perspective taken in this paper is in line with low-rank matrix completion approach, albeit not one that employs the trace heuristic seen in [28, 6, 92]. This work is more in line with non-convex approaches based on optimizing over a Riemannian manifold [32, 93], and extends the Riemannian approach of [54] to the EDMC basis case.
A recent work in [30] adopts a similar approach to us and considers solving the EDMC problem through Riemannian methods as well. In this work, the authors use a Riemannian conjugate method paired with an inexact line search method to minimize the following s-stress objective function:
@@ 7.2.1 Related Geometric Approaches to EDMC
where $g$ is the map defined by (2), $\bm{W}$ is a weight matrix to model noisy entries, and $\odot$ is the Hadamard product, and $\mathcal{P}_{\Omega}$ is defined as in (6). The analysis in [30] centers around the minimization of the s-stress function in (29) using a generalization of a Hager-Zhang line search method to a Riemannian quotient manifold. The main result in this work is that there exists an attractive basin for (29) that, with high probability, gives linear convergence to the ground truth provided an initialization in the basin. This result requires a Bernoulli sample complexity $p>C\frac{(\nu r)^{3}\log{n}}{n}$ , where $\nu$ is the incoherence of the ground truth matrix and $r$ is the rank. In contrast, our method also shows linear convergence in a local neighborhood and describes a strong initialization candidate for the noiseless EDMC recovery problem with provable high probability guarantees. We also provide robustness analysis for an EDMC problem perturbed by noise, and provide provable guarantees as well.
@@ 8 Numerical Results
All of the following experiments were conducted in MATLAB. The code used for the following experiments can be found in the GitHub repository at https://github.com/chandlersmith2/NonConvexEDMC.
@@ 8.1 Synthetic Data Experiment
Sphere (3D, $n=1002$) $3.38\times 10^{-7}$ $4.61\times 10^{-7}$ $6.12\times 10^{-7}$ $1.48\times 10^{-6}$ $8.40\times 10^{-3}$ $6.81\times 10^{-1}$
Cow (3D, $n=2601$) $4.41\times 10^{-7}$ $5.24\times 10^{-7}$ $6.04\times 10^{-7}$ $9.14\times 10^{-7}$ $2.47\times 10^{-4}$ $5.71\times 10^{-3}$
Swiss Roll (3D, $n=2048$) $3.85\times 10^{-7}$ $4.70\times 10^{-7}$ $5.81\times 10^{-7}$ $9.47\times 10^{-7}$ $1.56\times 10^{-6}$ $6.40\times 10^{-2}$
We note that the recovery completely fails for the sphere at $1\%$ sampling, while recovery is partially successful for the other two datasets. This is because the other datasets are larger while maintaining the same rank, allowing for better scaling in the low sampling regime. In Figure 4, we show an image of the reconstruction of the figures described in Table 2.
Figure 4: Reconstruction of the synthetic datasets referenced in Table 2. From left to right, the Bernoulli parameter is $0.03$, $0.02$, and $0.01$.
@@ 8.2 Comparison to Existing Methods
We provide an additional experiment to compare the efficacy of our algorithm DBRE to another provably convergent non-convex EDMC algorithm[34], which we refer as MatrixIRLS-EDMC . Let $r\in[2,10]$, and consider $n=100$ points sampled from $\mathrm{Unif}(S^{r-1})$, the uniform distribution on the sphere embedded in $r$ dimensions. As the number of degrees of freedom in a rank-$r$ $n\times n$ matrix is $nr-\frac{r(r-1)}{2}$, define the oversampling ratio $\rho$ as
@@
as $\mathbb{E}[|\Omega|]=pL$ for Bernoulli random sampling with parameter $p$. In Figure 5, we compare the oversampling ratio versus the dimension of the sphere in a transition plot. Black indicates complete failure, classified as a relative Gram matrix error larger than $10^{-3}$, and white indicates success. Each of these squares was run for 100 trials using DBRE and MatrixIRLS-EDMC. DBRE was initialized with 10 iterations of the Augmented Lagrangian algorithm in [28], and MatrixIRLS-EDMC was initialized using a least-squares methodology described in [34].
DBRE
MatrixIRLS-EDMC
Figure 5: Oversampling ratio $\rho$ versus dimension $r$ for $100$ points on the uniform distribution on $S^{r-1}$. Each parameter was tested 100 times.
As these experiments indicate, performance between MatrixIRLS for EDMC and Algorithm 1 is comparable, with Algorithm 1 performing slightly better overall, but most noticeably in the higher rank regime.
@@ 8.3 Experiments on Noisy Distance Measurements
Figure 6: Oversampling ratio $\rho$ versus noise level $10^{\gamma}$ for $100$ points drawn i.i.d. from $\mathrm{Unif}(S^{2})$. Each parameter was tested 500 times.
@@ 9 Conclusion and Future Work
In this work, we proposed a novel Riemannian gradient descent approach for solving the EDMC problem using a matrix completion approach on the manifold of rank-$r$ matrices in Algorithm 1. In a local neighborhood, we proved that Algorithm 1 exhibits linear convergence with high probability. To the authors’ knowledge, this is the first work to provide initialization guarantees for a non-convex approach to the EDMC problem. The convergence analysis of Algorithm 1 was predicated on a statistical understanding of coupled terms in a random operator, and required novel analysis to the matrix completion literature to our knowledge. For our method, we provided numerical results to underline its efficacy, and Algorithm 1 performs comparably to other state-of-the-art non-convex methods. Additionally, we provided robustness analysis and corresponding convergence guarantees. Finally, we provided a novel interpretation of incoherence in the EDMC setting, highlighting potential areas for development of non-uniform sampling methods in this field. This is a primary avenue of future interest, as improving the sample complexity through geometrically-optimal sampling schemes would represent a noteworthy development in the EDMC literature.
Acknowledgments
Abiy Tasissa and Chandler Smith acknowledge partial support from the National Science Foundation through grant DMS-2208392. HanQin Cai acknowledges partial support from the National Science Foundation through grant DMS-2304489. HanQin Cai and Abiy Tasissa appreciate insightful discussions with Prof. Rongjie Lai and Prof. Jian-Feng Cai at the initial stages of this work.
@@ Appendix A Properties of the Dual Bases and Concentration Inequalities
This section of the appendix details technical results about the specific dual bases, $\{\bm{w}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$ and $\{\bm{v}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$. These are needed to prove various technical lemmas throughout the work, but are particularly important in the proof of 5.3. Additionally, we provide the non-commutative and scalar Bernstein inequalities, as well as the Hanson-Wright inequality and the Davis-Kahan Theorem, all of which are leveraged throughout this work.
@@
$\mathbb{P}\left(\left\|\sum_{i=1}^{m}\bm{X}_{i}\right\|>t\right)\leq 2n\exp\left(-\frac{3t^{2}}{8\sigma^{2}}\right);$ (30)
and if $t>\frac{\sigma^{2}}{c}$,
@@
$\delta=\inf\left\{\left|\lambda-\hat{\lambda}\right|:\lambda\in[\lambda_{s},\lambda_{r}],\hat{\lambda}\in\left(-\infty,\hat{\lambda}_{s+1}\right)\cup\left(\hat{\lambda}_{r-1},\infty\right)\right\},$
@@
In particular, for rank-$r$ matrices $\bm{X},\hat{\bm{\bm{X}}}\succeq\bm{0}$ with eigenvectors corresponding to non-zero eigenvalues forming the columns of $\bm{V},\hat{\bm{V}}$, $\delta=\lambda_{r}$ and
@@
ne result that will be used throughout this work is a technique for constructing eigenvalue bounds through a vectorization technique. This result is as follows.
@@
[…]
Let $\tilde{\bm{H}}=[\langle\mathcal{P}_{U}\bm{w}_{\bm{\alpha}},\mathcal{P}_{U}\bm{w}_{\bm{\beta}}\rangle]\in\mathbb{R}^{L\times L}$, where $U$ is the row/column space of the true solution $\bm{X}=\bm{U}\bm{D}\bm{U}^{\top}$, which is rank-$r$, and where $\mathcal{P}_{U}$ is the projection operator onto $U$. It follows that
@@
as $\bm{w}_{\bm{\alpha}}\bm{w}_{\bm{\beta}}=\bm{w}_{\bm{\beta}}\bm{w}_{\bm{\alpha}}=\bm{0}$, where $\bm{0}$ is the zero matrix. Thus $\tilde{\bm{H}}$ is sparse, with each row having at most 2n32n-3 non-zero entries. The result follows from a Gershgorin argument and the entrywise bound derived from the incoherence condition above. ∎
@@
[…]
@@
The second statement follows from A.5 and the fact that $\mathcal{P}_{U}$ is an orthogonal projection operator. This concludes the proof. ∎
@@
Let $\bm{X}_{l}=\bm{U}_{l}\bm{D}_{l}\bm{U}_{l}^{\top}$ be a rank-$r$ matrix and $\mathbb{T}_{l}$ be the tangent space of $\mathcal{N}_{r}$ at $\bm{X}_{l}$. Let $\bm{X}=\bm{U}\bm{D}\bm{U}^{\top}$ be another rank-$r$ matrix, and $\mathbb{T}$ be the corresponding tangent space. Then
$\|\bm{U}_{l}\bm{U}_{l}^{\top}-\bm{U}\bm{U}^{\top}\|\leq\frac{\|\bm{X}_{l}-\bm{X}\|_{\mathrm{F}}}{\sigma_{r}(\bm{X})},\qquad\qquad\|\bm{U}_{l}\bm{U}_{l}^{\top}-\bm{U}\bm{U}^{\top}\|_{\mathrm{F}}\leq\frac{\sqrt{2}\|\bm{X}_{l}-\bm{X}\|_{\mathrm{F}}}{\sigma_{r}(\bm{X})},$
@@ Appendix B Restricted Isometry Results
$(\mathcal{S}_{\Omega}(\bm{X}))_{\bm{\alpha}}$ $=\begin{cases}\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle&\bm{\alpha}\in\Omega\\ 0&\bm{\alpha}\not\in\Omega\end{cases}$ (32)
@@
where $\{\xi_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$ are i.i.d. Bernoulli random variables that are $1$ with probability $p$ and 0 with probability 1p1-p.
@@
$(H^{-1}_{\mathrm{offdiag}})_{\bm{\alpha}\bm{\beta}}=\begin{cases}H^{\alpha\beta}&\bm{\alpha}\neq\bm{\beta};\\ 0&\bm{\alpha}=\bm{\beta}.\end{cases}$ (34)
@@
$\leq\exp\left(\lambda^{2}p\|\bm{A}\|_{\mathrm{F}}^{2}\|\bm{B}\|_{\mathrm{F}}^{2}\right),$
@@
Let $\hat{\mathcal{F}_{\Omega}}$ be either $\mathcal{P}_{\mathbb{T}}\mathcal{F}_{\Omega}\mathcal{P}_{\mathbb{T}}$ or $\mathcal{F}_{\Omega}$, and let $\hat{\mathcal{F}}_{\mathbb{I}}$ be $\mathcal{P}_{\mathbb{T}}\mathcal{F}_{\mathbb{I}}\mathcal{P}_{\mathbb{T}}$ or $\mathcal{F}_{\mathbb{I}}$, respectively. Let $\hat{\bm{w}}_{\bm{\alpha}}$ be either $\mathcal{P}_{\mathbb{T}}\bm{w}_{\bm{\alpha}}$ or $\bm{w}_{\bm{\alpha}}$, respectively. Let $\hat{\bm{H}}=[\langle\hat{\bm{w}}_{\bm{\alpha}},\hat{\bm{w}}_{\bm{\beta}}\rangle]\in\mathbb{R}^{L\times L}$ be the corresponding correlation matrix. For a ground truth rank $r$ $\nu$-incoherent matrix $\bm{X}$ with tangent space $\mathbb{T}$ on $\mathcal{N}_{r}$, we have that for any $\beta>1$, and with probability at least $1-2n^{1-\beta}$, and for $p\geq\frac{4\beta\log n}{3n}$, that
$\|\hat{\mathcal{F}}_{\Omega}-p\hat{\mathcal{F}}_{\mathbb{I}}\|\leq p\sqrt{\frac{8\beta\left(\max_{\bm{\alpha}}\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\right)\lambda_{\max}(\hat{\bm{H}})\log n}{3p}}.$
@@
This proof will follow a standard Bernstein argument. In order to make this argument, we need to build out $\hat{\mathcal{F}_{\Omega}}$ as a sum of random operators, we need to bound each, and then to bound the variance term. First, let $\{\xi_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$ be i.i.d. Bernoulli random variables that are 1 with probability $p$ and 0 with probability 1p1-p. It follows then that
@@
where the last inequality follows from 5.1. Next, we seek to bound the variance term, $\sigma^{2}=\left\|\sum_{\bm{\alpha}\in\mathbb{I}}\mathbb{E}\left[\bm{S}_{\bm{\alpha}}^{2}\right]\right\|$. To see this, first notice that
@@
[…]
@@
For a ground truth rank-$r$, $\nu$-incoherent matrix $\bm{X}$ with tangent space $\mathbb{T}$ on $\mathcal{N}_{r}$, we have that for any $\beta>1$ that if $p\geq\frac{4}{3}\frac{\beta\log{n}}{n}$, with probability at least $1-2n^{1-\beta}$
@@
We will prove this result using A.1. First, notice that
@@
where the first inequality follows from dropping the negative term, and the third inequality follows from 5.1 and $\|\bm{w}_{\bm{\alpha}}\|_{\mathrm{F}}=2$. Next, we note that
@@
where the first inequality follows from 5.1, the second inequality comes from A.5, and the final line comes from A.8. Next, notice that
@@
where the first inequality comes from $\|\bm{w}_{\bm{\alpha}}\|_{F}^{2}=4$, the second inequality comes from A.5, and the final line comes from $\lambda_{\max}(\tilde{\bm{H}})\leq\nu r$ in A.6. Taking $\sigma^{2}=\max\{\sigma_{1}^{2},\sigma_{2}^{2}\}=4p\nu r$, we get that for $t=\sqrt{\frac{32p\nu r\beta\log{n}}{3}}$ and $p\geq\frac{4}{3}\frac{\beta\log{n}}{{n}}$,
@@
[…]